Circular No. 6056

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994X IN ANONYMOUS GALAXY R. H. McNaught, Anglo-Australian Observatory, reports his discovery

of an apparent supernova of mag 18.5 on an R plate taken by M. Hartley with the U.K. Schmidt Telescope (UKST) on Aug. 15. The supernova is at α = 0^{h}15^{m}20^{s}.62, δ = −24^{o}53′50′′.5 (equinox 2000.0) and is located in a spiral arm, 1′′.0 east and 16′′.4 south of the galaxy’s center. The new object was confirmed on a CCD image taken by McNaught and G. J. Garradd with the 1.0-m telescope at Siding Spring on Aug. 16. No image appeared at this location on various UKST plates taken > 5 years ago. A nearby star of mag 17.5 has position end figures 21^{s}.70, 42′′.2.

X-RAY NOVA IN SCORPIUS C. A. Wilson, B. A. Harmon, S. N. Zhang, W. S. Paciesas, G. J. Fish-

man, M. Scott, B. C. Rubin, and R. B. Wilson report for the Compton Gamma Ray Observatory BATSE Team: “From an examination of individual occultations of GRO J1655–40, we find that the initial rise of the source intensity (∼ 10–90 percent) in the band 20–100 keV was between July 27.75 and 28.25 UT. The source was at the 0.85-Crab level from July 28.25 to ∼ Aug. 8 and then decreased to 0.66 Crab from Aug. 9 to 12 (band 20–100 keV; 10 percent uncertainty), with significant daily variations. Since Aug. 12, GRO J1655–40 has dropped more rapidly: 0.3 (± 0.03) Crab on Aug. 13 and 0.15 (± 0.03) Crab on Aug. 14–15. The spectrum extends to at least 200 keV in the early part of the outburst, then to at least 400 keV after Aug. 6. The spectrum softens during initial rise, from −2.5 to ∼ −3.0, and varies between −2.5 and −3.5 after July 28, similar to black-hole candidate x-ray novae. Power density spectra inspected through Aug. 5 show no low-frequency noise (10–500 mHz) in the hard x-ray region, which is unusual for such objects. Using data collected since Aug. 5, we obtain the following improved location: α = 16^{h}54^{m}.6, δ = −39^{o}54′ (equinox 2000.0; 1σ error along the southeast-to-northwest direction is 0^{o}.2 and along southwest-northeast is 0^{o}.1). This location is ∼ 0^{o}.13 from the optical candidate reported by Bailyn et al. (IAUC 6050).”

UZ BOOTIS This WZ-Sge-type variable may be in outburst for the first time in

16 years. Visual magnitude estimates: Aug. 16.458 UT, [15.1 (M. Iida, Nagano, Japan); 17.167, 13.9 (D. York, Abiquiu, NM).

1994 August 17 Daniel W. E. Green